A rectangular room is 6 meters longer than it is wide, and it’s perimeter is 24 meters. Find the dimension of the room.
A.) write an equation in terms of width, w that models the situation. Simplify it but do not solve.

Answer by ikleyn(52894)

.

Let the width be x meters;  then the length is (x+6) meters.


The perimeter is the sum of four lengths


    P = x + (x+6) + x + (x+6) = 2x + 2*(x+6),   or


    24 = 2x + 2*(x+6).


You may write it in this form


    x + (x+6) = 24/2

    
    x + (x+6) = 12.

--------------

I STOP at this point, since the assignment does nor requires to solve the equation.

The person, familiar with the subject, who just solved such problems many times, writes the equation

            x + (x+6) = half of the perimeter

without any explanations, assuming that everything is just crystally clear.